OpenStudy (llort):

Find n(12): 1, 2, 4, 7, 12, 20, 33, 54, 88

OpenStudy (amistre64):

the differences are fibonaccied

OpenStudy (anonymous):

terms in crease by 1 2 3 5 8 13 21 34 so next increments 55 89 144 the next one is the sum of previous two so n 12 = 88 + 55 + 89 + 144 = 376

OpenStudy (anonymous):

subtracting each by predecessor gives you this sequence: 1, 2, 3, 5, 8, 13, 21, 34... which is the Febunacci sequence ... 55, 89, 144, 223, 377 then just use this sequence to add each time to the next number... 88+55=143 143+89=232 232+144=376 so the n(12)=376